# Earliest Galaxy Evolution in the CANUCS+Technicolor fields: Galaxy Properties at $z\sim 10-16$ seen with the Full NIRCam Medium and Broad Band Filters

### v1

@@ Abstract
journal: ApJfacilities: HST/ACS, JWST/NIRCamsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), EAzY (G. B. Brammer et al., 2008), photutils (L. Bradley et al., 2023), Source Extractor (E. Bertin & S. Arnouts, 1996), emcee (D. Foreman-Mackey et al., 2013)
show]asada@kusastro.kyoto-u.ac.jp
@@ 1 Introduction
Finding and characterizing the earliest galaxies is one of the major quests in modern astronomy. The James Webb Space Telescope (JWST) has made a breakthrough in this field with its capability of finding $z\gtrsim 10$ galaxies. Numerous $z>10$ galaxy candidates have been identified with NIRCam imaging (e.g., Y. Harikane et al., 2023; P. G. Pérez-González et al., 2023; N. J. Adams et al., 2024; K. N. Hainline et al., 2024; C. J. Willott et al., 2024; M. Castellano et al., 2025; P. G. Pérez-González et al., 2025). The statistical search for high-$z$ galaxy candidates with NIRCam imaging is in particular essential to probe the earliest galaxy evolution. Tracking the early evolution of galaxy number density, e.g., the UV luminosity function (LF), directly unveils how first galaxies have emerged from the initial matter perturbation and how they have grown with baryon accretion in dark matter halos (e.g., R. J. Bouwens et al., 2021; Y. Harikane et al., 2022).
@@
Various NIRCam imaging surveys have been carried out in the first three years of JWST observation, most of which employed the standard Broad-band (BB) filter configurations. This has been a good strategy as BB filters are wide in wavelength coverage, offering good spectral energy distribution (SED) coverage but also facilitating deep photometry with efficient integration times. Given that the Lyman break feature is strong and does not require high spectral resolution, the BB approach in the initial years of JWST has been sensible and highly efficient at discovering $z>10$ galaxy candidates. Thus far, many $z_{\rm phot}\gtrsim 10$ galaxy samples are built relying on the BB-only photometry data, including galaxy candidates even up to $z_{\rm phot}\sim 25$ (e.g., M. Castellano et al., 2025; P. G. Pérez-González et al., 2025). However, the selection of high-$z$ galaxies with BB-only data needs to be treated with caution, because the BB photometry spectral energy distributions (SEDs) of interlopers at lower-$z$ with strong emission lines combined with a Balmer break or a red, dusty continuum can mimic the high-$z$ galaxy SED and contaminate the high-$z$ galaxy sample. Perhaps the most prominent example of this is the galaxy CEERS-9331, which was originally identified as a bright $z\sim 16$ galaxy candidate by several independent authors based on the seven BB filters + F410M photometric data (e.g., C. T. Donnan et al., 2023; Y. Harikane et al., 2023), but subsequent NIRSpec observation confirmed it as a dusty extreme emission line galaxy at $z=4.9$ (P. Arrabal Haro et al., 2023).
@@
Particularly at $z>9.5$, all strong rest-frame optical emission lines are redshifted beyond NIRCam filter coverage, and the Lyman-$\alpha$ break at $\lambda_{\rm rest}=1216\ {\rm\AA}$ is the only signature to identify $z>9.5$ galaxy candidates from photometry. The combination of NIRCam Medium-band (MB) and BB filter observations is one way to minimize contamination when building robust $z>9.5$ galaxy samples (e.g., N. J. Adams et al., 2025). NIRCam MB observations have successfully shown that they can identify strong emission lines from the flux excess in MB filters (e.g., S. Withers et al., 2023; G. T. E. Sarrouh et al., 2024; K. A. Suess et al., 2024; N. Martis et al., 2025) and therefore they are extremely efficient at identifying low-$z$ strong line emitters that contaminate high-$z$ candidates such as in P. Arrabal Haro et al. (2023). MB+BB observations also allow us to precisely locate the sharp break in the continuum at the Lyman break. Thus MB+BB filter observations should unambiguously distinguish low-$z$ interlopers with strong emission lines vs real high-$z$ galaxies with sharp continuum drop-outs.
@@
Not only can MBs efficiently rule out interlopers, they can also help properly classify real candidates previously considered more likely to be low-$z$ contaminants with BB-only data. One good showcase of this is GS-z14-0 at $z_{\rm spec}=14.18$ (S. Carniani et al., 2025). GS-z14-0 was originally identified from NIRCam BB imaging observations in Cycle 1, but it was thought to be at $z\sim 3.5$ with the BB-only data (C. C. Williams et al., 2024). After obtaining MB observations in Cycle 2, GS-z14-0 was identified as a secure $z\sim 14$ galaxy candidate (B. Robertson et al., 2024), and the subsequent spectroscopic follow-up observation proved it to be at $z_{\rm spec}=14.18$ (S. Carniani et al., 2024, 2025).
@@
Specifically, the combination of F140M, F150W, and F162M is powerful in selecting $z\sim 10-13$ galaxies, and that of F182M, F200W, and F210M in selecting $z\sim 13-16$ galaxies, as they simultaneously cover the Lyman break wavelength range with two MBs and one BB overlapping each other, and they can locate the sharp drop-out from photometry (or strong emission line excess for low-$z$ interlopers). B. Robertson et al. (2024) exploited a deep MB+BB NIRCam survey program, the JADES Origins Field (JOF; D. J. Eisenstein et al., 2023), to derive the galaxy UV LF at $z>9.5$ based on 14 NIRCam filter images (seven MBs and seven BBs) in a very deep, single NIRCam pointing. As pointed out by B. Robertson et al. (2024), galaxy abundance measurements in such a small area are affected by cosmic variance, and it is necessary to enlarge the survey volume for $z>10$ galaxies with a comprehensive set of MB+BB filters.
@@
In this paper, we present UV LF measurements at $z>9.5$ using data from the CAnadian NIRISS Unbiased Cluster Survey (CANUCS; GTO-1208, PI: Willott; C. J. Willott et al., 2022) in Cycle 1 and the JWST in Technicolor program (TEC; GO-3362, PI: Muzzin) in Cycle 2. Combining CANUCS+TEC provides the full set of NIRCam BB+MB filter observations (+ two narrow-band filters) in three independent NIRCam pointings. Observing three independent sight lines with the rich MB+BB filters adds an invaluable constraint on the early UV LF evolution, by building a new, very clean $z>10$ galaxy sample in addition to the previous work in the JOF. We also explore the potential bias due to low-$z$ interloper contamination in BB-only data selection of $z\gtrsim 10$ galaxies, by comparing our main $z\gtrsim 10$ candidates built from MB+BB observations and those from a degraded BB-only photometry catalog.
The paper is structured as follows. Sec. 2 describes the imaging data and data processing we use in this work. Sec. 3 presents the sample selection of high-$z$ galaxies and the completeness simulation of our selection. Sec. 4 gives the main result of this paper, including the UV LF at $z>9.5$ and the redshift evolution of the UV luminosity density. Sec. 5 is the discussion focusing on what can affect measurements of high-$z$ galaxy abundance in JWST surveys, including cosmic variance and low-$z$ interlopers in BB-only data selections. The section also discusses the potential bias by inclusion of low-$z$ interlopers with BB-only data in various aspects of galaxy properties at $z>10$ and future prospects from scheduled JWST Cycle 4 observations. Sec. 6 finally provides the summary of the paper. Throughout the paper, we assume a flat $\Lambda$CDM cosmology with $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$, and $H_{0}=70\ {\rm km\ s^{-1}\ Mpc^{-1}}$, and all magnitudes are quoted in the AB system (J. B. Oke & J. E. Gunn, 1983).
@@ 2.1 Imaging Data and Photometry
We use JWST/NIRCam imaging data from CANUCS in Cycle 1 and TEC in Cycle 2 of the three NIRCam flanking fields (NCFs). The overall survey description and full details of the imaging reduction are presented in G. T. E. Sarrouh et al. (2025, SA25 hereafter). Combining the multi-epoch two surveys, a total of three NIRCam pointings are observed with all NIRCam BB and MB filters111except for the MACS 1149 flanking field: F162M and F250M are missing due to a program definition error, in addition to two narrow-band filters of F164N and F187N. Each NIRCam pointing is an associated flanking field of the strong lensing cluster field (MACS 0416, MACS 1149, and A 370), and we refer to these flanking fields as NCF fields. As we only use the flanking fields, the effect of gravitational lensing is negligibly small in this work. Table 1 presents the limiting magnitudes of all available NIRCam filters in each field. The table also includes the information about the observing cycle (Cycle 1 by CANUCS program or Cycle 2 by TEC program).
Table 1: Summary of JWST/NIRCam and HST/ACS observations in CANUCS+Technicolor fields.
@@
footnotetext: Notes. For each filter, the 3-sigma limiting magnitudes (in ABmag) are quoted. The limiting magnitudes are measured in $0.\!\!^{\prime\prime}16$-diameter aperture, by randomly distributing empty apertures in the background sky region of each image (without PSF-convolution) and performing aperture correction based on the PSF EEs.
We also utilize HST optical imaging observations in F435W, F606W, and F814W filters on Advanced Camera for Surveys (ACS), taken in previous observations by the Hubble Frontier Fields Parallels program (J. M. Lotz et al., 2017). The HST/ACS footprints partially overlap with the NIRCam area (typically $\sim 50\ \%$ of NIRCam field-of-views), so we use HST/ACS photometry when available for photometric redshift computation (Sec. 2.2), but do not adopt S/N cuts in HST/ACS filter images in the sample selection. There are also several HST Wide Field Camera 3 infra-red filter images available in these fields, but we do not use them in this work, given the wavelength overlap with NIRCam and the superior quality of NIRCam images in resolution and sensitivity.
We exploit the background-subtracted images and the source detection catalogs given by SA25, but we perform custom photometry to maximize the signal-to-noise ratio (S/N) for faint, small sources barely resolved with NIRCam imaging. Namely, at the position of each detected source, we perform fixed aperture photometry with a $0.\!\!^{\prime\prime}16$-diameter aperture on the background-subtracted images that have not been convolved to a common PSF. We correct the aperture photometry in each filter based on the PSF encircled energies (EEs) of the filter, since the galaxies are small. The photometry is then corrected for Galactic extinction with the E. L. Fitzpatrick (1999) extinction curve assuming $R_{V}=3.1$, to obtain the color (i.e., SEDs) of each detected source. The total flux of each source is estimated by scaling the SED with the Kron-to-aperture-corrected flux ratios in the F277W image.
@@ 3 Sample Selection
Table 2: High-redshift galaxy candidates in CANUCS+Technicolor fields.
@@
CANUCS-5223684 177.414290 22.32748 $9.65^{+0.32}_{-0.35}$ $-18.37$ $-2.14\pm 0.22$ $17.1\pm 2.3$ 4.96e-04
CANUCS-3209488 64.169709 -24.10793 $9.69^{+0.24}_{-0.36}$ $-19.23$ $-1.88\pm 0.27$ $487.9\pm 27.3$aaThe source appears to have two components in the NIRCam image, but fails to fit with two Sérsic components and results of fitting a single component are quoted. 2.76e-09
CANUCS-2215881 40.056129 -1.61568 $9.78^{+0.13}_{-0.43}$ $-19.09$ $-2.20\pm 0.14$ $109.8\pm 83.3$ 8.00e-11
CANUCS-5221204 177.386651 22.30266 $10.15^{+0.02}_{-1.42}$ $-18.89$ $-2.67\pm 0.37$ $207.6\pm 91.1$ 2.17e-03
CANUCS-2219868 40.032579 -1.60055 $10.32^{+0.15}_{-0.26}$ $-19.52$ $-2.33\pm 0.16$ $112.4\pm 32.6$bbFitted with two Sérsic components, and the size of the main clump is quoted. 1.66e-04
CANUCS-5208401 177.404008 22.32253 $10.69^{+0.33}_{-0.38}$ $-18.75$ $-1.93\pm 0.30$ $15.9\pm 2.1$ 6.52e-02
CANUCS-5203757 177.408240 22.28612 $11.23^{+0.14}_{-0.10}$ $-19.77$ $-2.18\pm 0.07$ $15.3\pm 2.0$ 9.92e-19
CANUCS-3214552 64.133557 -24.07987 $15.43^{+0.26}_{-0.57}$ $-18.98$ $-1.80\pm 0.34$ $29.2\pm 1.3$ 5.05e-03
footnotetext: (1) Source IDs in the CANUCS catalog by SA25. (2) Right Ascension in J2000. (3) Declination in J2000. (4) Best photo-$z$ estimations. Uncertainties are from the 16th- to 84th-percentile range of $P(z)$. (5) Absolute rest UV magnitudes. (6) UV slopes measured from NIRCam photometry. (7) The intrinsic, PSF-deconvolved effective radii $r_{\rm eff}$ in the NIRCam F277W image. (8) Low-$z$ solution probabilities.
@@
where $S/N_{\rm Blue}<3$ means the signal-to-noise ratio (S/N) in bluer filters (F070W, F090W, and F115W) are all less than 3. Requiring S/Ns smaller than 3 in three independent filters would remove $\sim 1\ \%$ of real sources by the random errors. If we required S/Ns smaller than 2 instead, we would miss $\sim 14\ \%$ of real candidates due to the random noise. The limit of low-$z$ solution probability is applied to select only high-$z$ sources with high confidence level and remove possible low-$z$ interlopers that have double-peaked $P(z)$. We also adopt the redshift upper limit of $16$ to remove any spurious NIRCam long wave (LW)-only sources: this is because the relatively large wavelength coverage gap between NIRCam short-wave (SW) filters and LW filters, unmasked bad pixels in the LW detectors, broader PSF size at longer wavelength, or the combination of them can easily produce fake high-$z$ candidates that are detected only in NIRCam LW filters. The motivation of this work is to provide the cleanest sample of $z>9.5$ galaxies possible, relying on the rich NIRCam MB+BB observations that are capable to robustly capture the Lyman break, thus we limit the redshift up to $z<16$ where the dropout can still be detected with F200W and F210M on the SW detectors.

### v2

@@ Abstract
journal: ApJfacilities: HST/ACS, JWST/NIRCamsoftware: astropy (2013A&A...558A..33A; 2018AJ....156..123A; 2022ApJ...935..167A), EAzY (Brammer2008ApJ), photutils (larry_bradley_2023_7946442), Source Extractor (1996A&AS..117..393B), emcee (emcee)
show]yoshi.asada@utoronto.ca
@@ 1 Introduction
Finding and characterizing the earliest galaxies is one of the major quests in modern astronomy. The James Webb Space Telescope (JWST) has made a breakthrough in this field with its capability of finding $z\gtrsim 10$ galaxies. Numerous $z>10$ galaxy candidates have been identified with NIRCam imaging (e.g., Naidu2022ApJ; Bouwens2023MNRAS; Donnan2023; Donnan2023MNRAS; Harikane2023; Perez_Gonzalez2023; Adams2024ApJ; Hainline2024; Willott2024ApJ; Castellano2025arXiv; Perez_Gonzalez2025). The statistical search for high-$z$ galaxy candidates with NIRCam imaging is in particular essential to probe the earliest galaxy evolution. Tracking the early evolution of galaxy number density, e.g., the UV luminosity function (LF), directly unveils how first galaxies have emerged from the initial matter perturbation and how they have grown with baryon accretion in dark matter halos (e.g., Bouwens2021; Harikane2022).
@@
Various NIRCam imaging surveys have been carried out in the first three years of JWST observation, most of which employed the standard Broad-band (BB) filter configurations. This has been a good strategy as BB filters are wide in wavelength coverage, offering good spectral energy distribution (SED) coverage but also facilitating deep photometry with efficient integration times. Given that the Lyman break feature is strong and does not require high spectral resolution, the BB approach in the initial years of JWST has been sensible and highly efficient at discovering $z>10$ galaxy candidates. Thus far, many $z_{\rm phot}\gtrsim 10$ galaxy samples are built relying on the BB-only photometry data, including galaxy candidates even up to $z_{\rm phot}\sim 25$ (e.g., Castellano2025arXiv; Perez_Gonzalez2025). However, the selection of high-$z$ galaxies with BB-only data needs to be treated with caution, because the BB photometry spectral energy distributions (SEDs) of interlopers at lower-$z$ with strong emission lines combined with a Balmer break or a red, dusty continuum can mimic the high-$z$ galaxy SED and contaminate the high-$z$ galaxy sample. Perhaps the most prominent example of this is the galaxy CEERS-9331, which was originally identified as a bright $z\sim 16$ galaxy candidate by several independent authors based on the seven BB filters + F410M photometric data (e.g., Donnan2023; Harikane2023), but subsequent NIRSpec observation confirmed it as a dusty extreme emission line galaxy at $z=4.9$ (Arrabal_Haro2023).
@@
Particularly at $z>9.5$, all strong rest-frame optical emission lines are redshifted beyond NIRCam filter coverage, and the Lyman-$\alpha$ break at $\lambda_{\rm rest}=1216\ {\rm\AA }$ is the only signature to identify $z>9.5$ galaxy candidates from photometry. The combination of NIRCam Medium-band (MB) and BB filter observations is one way to minimize contamination when building robust $z>9.5$ galaxy samples (e.g., Adams2025). NIRCam MB observations have successfully shown that they can identify strong emission lines from the flux excess in MB filters (e.g., Withers2023; Sarrouh2024ApJ; Suess2024ApJ; Martis2025arXiv) and therefore they are extremely efficient at identifying low-$z$ strong line emitters that contaminate high-$z$ candidates such as in Arrabal_Haro2023. MB+BB observations also allow us to precisely locate the sharp break in the continuum at the Lyman break. Thus MB+BB filter observations should unambiguously distinguish low-$z$ interlopers with strong emission lines vs real high-$z$ galaxies with sharp continuum drop-outs.
@@
Not only can MBs efficiently rule out interlopers, they can also help properly classify real candidates previously considered more likely to be low-$z$ contaminants with BB-only data. One good showcase of this is GS-z14-0 at $z_{\rm spec}=14.18$ (Carniani2025AAS). GS-z14-0 was originally identified from NIRCam BB imaging observations in Cycle 1, but it was thought to be at $z\sim 3.5$ with the BB-only data (Williams2024ApJ). After obtaining MB observations in Cycle 2, GS-z14-0 was identified as a secure $z\sim 14$ galaxy candidate (Robertson2024ApJ), and the subsequent spectroscopic follow-up observation proved it to be at $z_{\rm spec}=14.18$ (Carniani2024; Carniani2025AAS).
@@
Specifically, the combination of F140M, F150W, and F162M is powerful in selecting $z\sim 10-13$ galaxies, and that of F182M, F200W, and F210M in selecting $z\sim 13-16$ galaxies, as they simultaneously cover the Lyman break wavelength range with two MBs and one BB overlapping each other, and they can locate the sharp drop-out from photometry (or strong emission line excess for low-$z$ interlopers; Figure 1). Robertson2024ApJ exploited a deep MB+BB NIRCam survey program, the JADES Origins Field (JOF; Eisenstein2023arXiv), to derive the galaxy UV LF at $z>9.5$ based on 14 NIRCam filter images (seven MBs and seven BBs) in a very deep, single NIRCam pointing. As pointed out by Robertson2024ApJ, galaxy abundance measurements in such a small area are affected by cosmic variance, and it is necessary to enlarge the survey volume for $z>10$ galaxies with a comprehensive set of MB+BB filters.
Figure 1: Illustration showing the NIRCam filter transmission curves and the Lyman-$\alpha$ break feature in $z>10$ galaxies. The MB filters are shown by filled curves.
@@
In this paper, we present UV LF measurements at $z>9.5$ using data from the CAnadian NIRISS Unbiased Cluster Survey (CANUCS; GTO-1208, PI: Willott; Willott2022PASP) in Cycle 1 and the JWST in Technicolor program (TEC; GO-3362, PI: Muzzin) in Cycle 2. Combining CANUCS+TEC provides the full set of NIRCam BB+MB filter observations (+ two narrow-band filters) in three independent NIRCam pointings. Observing three independent sight lines with the rich MB+BB filters adds an invaluable constraint on the early UV LF evolution, by building a new, very clean $z>10$ galaxy sample in addition to the previous work in the JOF. We also explore the potential bias due to low-$z$ interloper contamination in BB-only data selection of $z\gtrsim 10$ galaxies, by comparing our main $z\gtrsim 10$ candidates built from MB+BB observations and those from a degraded BB-only photometry catalog.
[…]
@@ 2.1 Imaging Data and Photometry
We use JWST/NIRCam imaging data from CANUCS in Cycle 1 and TEC in Cycle 2 of the three NIRCam flanking fields (NCFs). The overall survey description and full details of the imaging reduction are presented in Sarrouh2025arXiv. Combining the multi-epoch two surveys, a total of three NIRCam pointings are observed with all NIRCam BB and MB filters111except for the MACS 1149 flanking field: F162M and F250M are missing due to a program definition error, in addition to two narrow-band filters of F164N and F187N. Each NIRCam pointing is an associated flanking field of the strong lensing cluster field (MACS 0416, MACS 1149, and A 370), and we refer to these flanking fields as NCF fields. As we only use the flanking fields, the effect of gravitational lensing is negligibly small in this work. Table 1 presents the limiting magnitudes of all available NIRCam filters in each field. The table also includes the information about the observing cycle (Cycle 1 by CANUCS program or Cycle 2 by TEC program).
Table 1: Summary of JWST/NIRCam and HST/ACS observations in CANUCS+Technicolor fields. footnotetext: Notes. For each filter, the 3-sigma limiting magnitudes (in ABmag) are quoted. The limiting magnitudes are measured in $0.\!\!^{\prime\prime}16$-diameter aperture, by randomly distributing empty apertures in the background sky region of each image (without PSF-convolution) and performing aperture correction based on the PSF $E$s.
@@
We also utilize HST optical imaging observations in F435W, F606W, and F814W filters on Advanced Camera for Surveys (ACS), taken in previous observations by the Hubble Frontier Fields Parallels program (Lotz2017ApJ). The HST/ACS footprints partially overlap with the NIRCam area (typically $\sim 50\ \%$ of NIRCam field-of-views), so we use HST/ACS photometry when available for photometric redshift computation (Sec. 2.2), but do not adopt S/N cuts in HST/ACS filter images in the sample selection. There are also several HST Wide Field Camera 3 infra-red filter images available in these fields, but we do not use them in this work, given the wavelength overlap with NIRCam and the superior quality of NIRCam images in resolution and sensitivity.
We exploit the background-subtracted images and the source detection catalogs given by SA25, but we perform custom photometry to maximize the signal-to-noise ratio (S/N) for faint, small sources barely resolved with NIRCam imaging. In SA25, the source detection is performed on the $\chi_{\rm mean}$ image following Drlica-Wagner2018ApJS, which is built by co-adding all available JWST and HST-optical filters. At the position of each detected source, we perform fixed aperture photometry with a $0.\!\!^{\prime\prime}16$-diameter aperture on the background-subtracted images that have not been convolved to a common PSF. We correct the aperture photometry in each filter based on the PSF encircled energies (EEs) of the filter, since the galaxies are small. The photometry is then corrected for Galactic extinction with the Fitzpatrick1999PASP extinction curve assuming $R_{V}=3.1$, to obtain the color (i.e., SEDs) of each detected source. The total flux of each source is estimated by scaling the SED with the Kron-to-aperture-corrected flux ratios in the F277W image. The photometric errors are computed by measuring 2000 empty aperture fluxes in the background sky region and applying the same correction factors as the fixed aperture fluxes (i.e., color correction, Galactic dust attenuation correction, and aperture correction).
@@ 3 Sample Selection
Table 2: High-redshift galaxy candidates in CANUCS+Technicolor fields. footnotetext: (1) Source IDs in the CANUCS catalog by SA25. (2) Right Ascension in J2000. (3) Declination in J2000. (4) Best photo-$z$ estimations. Uncertainties are from the 16th- to 84th-percentile range of $P(z)$. (5) Absolute rest UV magnitudes. (6) UV slopes measured from NIRCam photometry. (7) The intrinsic, PSF-deconvolved effective radii $r_{\rm eff}$ in the NIRCam F277W image. (8) Low-$z$ solution probabilities.
@@
CANUCS-5223684 177.414290 22.32748 $9.65^{+0.32}_{-0.35}$ 18.37-18.37 $2.14\pm 0.22$ $17.1\pm 2.3$ 4.96e-04
CANUCS-3209488 64.169709 -24.10793 $9.69^{+0.24}_{-0.36}$ 19.23-19.23 $1.88\pm 0.27$ $487.9\pm 27.3$aaThe source appears to have two components in the NIRCam image, but fails to fit with two Sérsic components and results of fitting a single component are quoted. 2.76e-09
CANUCS-2215881 40.056129 -1.61568 $9.78^{+0.13}_{-0.43}$ 19.09-19.09 $2.20\pm 0.14$ $109.8\pm 83.3$ 8.00e-11
CANUCS-5221204 177.386651 22.30266 $10.15^{+0.02}_{-1.42}$ 18.89-18.89 $2.67\pm 0.37$ $207.6\pm 91.1$ 2.17e-03
CANUCS-2219868 40.032579 -1.60055 $10.32^{+0.15}_{-0.26}$ 19.52-19.52 $2.33\pm 0.16$ $112.4\pm 32.6$bbFitted with two Sérsic components, and the size of the main clump is quoted. 1.66e-04
CANUCS-5208401 177.404008 22.32253 $10.69^{+0.33}_{-0.38}$ 18.75-18.75 $1.93\pm 0.30$ $15.9\pm 2.1$ 6.52e-02
CANUCS-5203757 177.408240 22.28612 $11.23^{+0.14}_{-0.10}$ 19.77-19.77 $2.18\pm 0.07$ $15.3\pm 2.0$ 9.92e-19
CANUCS-3214552 64.133557 -24.07987 $15.43^{+0.26}_{-0.57}$ 18.98-18.98 $1.80\pm 0.34$ $29.2\pm 1.3$ 5.05e-03
@@
where $S/N_{\rm Blue}<3$ means the signal-to-noise ratio (S/N) in bluer filters (F070W, F090W, and F115W) are all less than 3. Requiring S/Ns smaller than 3 in three independent filters would remove $\sim 0.3\ \%$ of real sources by the random errors. If we required S/Ns smaller than 2 instead, we would miss $\sim 6\ \%$ of real candidates due to the random noise. The limit of low-$z$ solution probability is applied to select only high-$z$ sources with high confidence level and remove possible low-$z$ interlopers that have double-peaked $P(z)$. We also adopt the redshift upper limit of $16$ to remove any spurious NIRCam long wave (LW)-only sources: this is because the relatively large wavelength coverage gap between NIRCam short-wave (SW) filters and LW filters, unmasked bad pixels in the LW detectors, broader PSF size at longer wavelength, or the combination of them can easily produce fake high-$z$ candidates that are detected only in NIRCam LW filters. The motivation of this work is to provide the cleanest sample of $z>9.5$ galaxies possible, relying on the rich NIRCam MB+BB observations that are capable to robustly capture the Lyman break, thus we limit the redshift up to $z<16$ where the dropout can still be detected with F200W and F210M on the SW detectors.